OpenStudy (_t_b_w_):

Help? I will medal and fan :) olve for x: |3x + 12| = 18 x = 2, x = –10 x = 2, x = –2 x = 10, x = –10 x = –2, x = 10

Awolflover1 (awolflover1):

\[3x+12=18\]

Awolflover1 (awolflover1):

Subtract 12 from both sides \[3x+12−12=18−12\]

OpenStudy (_t_b_w_):

Ok

OpenStudy (_t_b_w_):

I got two as the answer. So would it be B?

Awolflover1 (awolflover1):

close but incorrect

Awolflover1 (awolflover1):

let's do the other one so you can get your answer

OpenStudy (_t_b_w_):

ok

Awolflover1 (awolflover1):

\[3x+12=−18\] \[3x+12−12=−18−12\]

Awolflover1 (awolflover1):

\[3x=−30\] Now Divide both sides by 3

OpenStudy (_t_b_w_):

Ok

OpenStudy (_t_b_w_):

I got -10. So it would be A?

Awolflover1 (awolflover1):

yep (:

OpenStudy (_t_b_w_):

thank you so much :)

Awolflover1 (awolflover1):

You're Welcome! ^_^

OpenStudy (_t_b_w_):

I just medaled you and will fan :)